OpenStudy (anonymous):

What is the probability of tossing heads on a coin and a number greater than 4 on a number cube? A. 1/6 B. 1/4 C. 1/3 D. 2/3

OpenStudy (anonymous):

I think its A

OpenStudy (anonymous):

P(head) = 1/2 P(x>4) = 2/6 P = (1/2)(2/6) = 1/6

OpenStudy (anonymous):

thanks guys